A particle of mass 10²  Kg is moving along the positive x-axis under the influence of a force F(x) = -k/2x²  where k= 10-²  NM² . at time t=0 it is at x=1.0 m and its velocity is v=0  (a) Find its velocity when it reaches x= 0.5 m

Grade:11

1 Answers

Shivam
9 years ago
Change in Kinetic energy = work done 
½ mv2 – 1/2 mu2 =∫ F.dx =>1/2 m(vf2-v2) = [k/x3]]
½ * 102 (vf2 )= 10-2/(1/.125 – 1/1)
vf2= 14 * 10-4
vf = 3.74 * 10-2 m/s